OpenStudy (anonymous):

What is the period of the sinusoidal voltage f(t)=50cos(4πt-30°)?

OpenStudy (anonymous):

0.5 secs

OpenStudy (anonymous):

w=2(pi)f f=2 T=1/f =0.5

OpenStudy (anonymous):

yaa its 0.5

OpenStudy (anonymous):

Thank you very much!!!!!!!

OpenStudy (anonymous):

.5

OpenStudy (anonymous):

T=2(pi)/(omega) or omega=4pi dont T=0,5

OpenStudy (anonymous):

Some how I remember answering this very same question yesterday. What is it you don't understand?